# A Computational Theory and Semi-Supervised Algorithm for Clustering

### v1

@@ 4 Conclusion
Secondly, a novel semi-supervised clustering algorithm is proposed whose kernel is based on the anomaly detection work by Mohammad [8]. This work gave a precise computable definition of what constitutes an anomaly and provides a compelling rationale for accepting such a definition. Additionally, an efficient algorithm to compute anomalies based on the Euclidean distance from the median is provided. The clustering algorithm operates on numerical data and utilises a small amount of labelled examples for initialisation, guidance, and subsequent expansion of clusters. The algorithm handles noisy labelled data by ejecting distinctly incorrect points from the groups. Instead of aiming for a complete partitioning of all the data, the algorithm focuses on clustering the data, thus ensuring that anomalies and unknown clusters are not forcibly assigned to existing clusters. Such instances are left to be analysed in subsequent steps.
The results of the semi-supervised clustering algorithm are compared with those of popular unsupervised clustering algorithms, namely $k$-means and DBSCAN. The synthetic data examples showcased the algorithm’s ability to build clusters even with a relatively small number of examples per cluster and accommodated clusters of different densities. In contrast, $k$-means and DBSCAN exhibited undesired results in these examples due to their inherent constraints and assumptions. In particular $k$-means gave unintuitive partitions that split clusters or assigned clearly anomalous points to a cluster. On the other hand, DBSCAN proved to be sensitive to the supplied parameters where clusters either got merged or split because of its inability to locally adapt.
The incorporation of dimension reduction techniques, particularly UMAP, in conjunction with semi-supervised clustering, has proven to be highly advantageous. This is exemplified through the analysis of image and text data examples. The employed workflow, which encompasses dimension reduction for visualisation, sample labelling, and clustering, offers significant benefits in the identification of clusters. Notably, it recognises the importance of not forcibly assigning every example, including anomalies and points from unknown clusters, to specific clusters. Instead, the approach allows for a more nuanced treatment of data points, facilitating the identification of clusters while acknowledging the presence of ungrouped instances. These examples also showed that user interaction and sampling remain important even after UMAP visualisation, as points appearing visually close together may not belong to the same group. An example illustrating this phenomenon can be found in the MNIST digits mapping (Figure 13), where the clusters of digits 8’s and 1’s exhibit visual proximity while representing separate clusters.
@@
Clustering is an interactive and iterative process that combines exploration and analysis. Merely applying an algorithm without understanding the data and the problem space is not recommended, and results are never accepted blindly. During the exploratory phase, it is important to gain familiarity with the data and establish certain assumptions that constrain solutions, such as the type and number of clusters, the possibility of clusters forming hyper-spherical groupings, algorithm parameters, or the labelling of points for seeding and guiding an algorithm. In the presented semi-supervised clustering algorithm, the explicit constraints are taken to be labelled examples as opposed to parameter specification, and implicit is the choice of grouping law and measure. However, note the difference between knowledge of a few labelled examples per cluster and estimating parameters such as the $k$ in $k$-means or $mps$ and $eps$ in DBSCAN. In the former case clusters can be gradually built over the labelled subset, and then expanded upon, without requiring knowledge of samples from all the clusters. Furthermore, being parameter-free relieves users of one of the main challenges commonly encountered in unsupervised learning problems.
@@
In future research, the exploration of additional grouping principles and measures, extending beyond proximity and the Euclidean distance to the median, will be undertaken. This expansion aims to enhance the versatility and effectiveness of the clustering algorithm. Moreover, the inherent speed and nature of the algorithm also makes it well-suited for online clustering scenarios. Consequently, an extension of the algorithm to handle the assignment of new points to existing clusters and the identification of anomalies that may potentially form new clusters will be investigated.

### v2

@@ 4 Conclusion
Secondly, a novel semi-supervised clustering algorithm is proposed whose kernel is based on the anomaly detection work by Mohammad [22]. This work gave a precise computable definition of what constitutes an anomaly and provides a compelling rationale for accepting such a definition. Additionally, an efficient algorithm to compute anomalies based on Expectation and Euclidean distance from the median is provided. The clustering algorithm operates on numerical data and utilises a small amount of labelled examples for initialisation, guidance, and subsequent expansion of clusters. Results have shown that as a rule of thumb, $10$-$30$ points per known cluster, give effective clustering results. The algorithm handles noisy labelled data by ejecting unexpected points (anomalies) from the groups. Instead of aiming for a complete partitioning of all the data, the algorithm focuses on clustering the data, thus ensuring that anomalies and unknown clusters are not forcibly assigned to existing clusters. If full partitioning is required, points can optionally be assigned to their nearest cluster based on anomaly scores. Such instances are left to be analysed in subsequent steps. Crucially, the proposed clustering algorithm is parameter-free, fast, and efficient—eliminating the need for users to manually set parameters.
The qualitative analysis over synthetic datasets and subsequent quantitative evaluation using established external validation metrics demonstrated the effectiveness of the proposed algorithm to consistently obtain top-tier or competitive performance. The results thus indicate its potential for real-world applications and as a novel contribution to the field of clustering. The use of dimensionality reduction techniques—particularly UMAP—in combination with clustering has also proven beneficial, as demonstrated in both image and text data analyses. The adopted workflow, which integrates dimensionality reduction for visualisation with sampling, labelling, and clustering, enhances cluster identification while also improving overall efficiency.
@@
Clustering is an interactive and iterative process that combines exploration and analysis. Merely applying an algorithm without understanding the data and the problem space is not recommended, and results are never accepted blindly. During the exploratory phase, it is important to gain familiarity with the data and establish certain assumptions that constrain solutions, such as the type and number of clusters, the possibility of clusters forming hyper-spherical groupings, algorithm parameters, or the labelling of points for seeding and guiding an algorithm. In the presented semi-supervised clustering algorithm, the explicit constraints are taken to be labelled examples as opposed to parameter specification, and implicit is the choice of grouping law and measure. However, note the difference between knowledge of a few labelled examples per cluster and estimating parameters such as the $k$ in $k$-means or $mps$ and $eps$ in DBSCAN. In the former case clusters can be gradually built over the labelled subset, and then expanded upon, without requiring knowledge of samples from all the clusters or requiring parameter estimation.
@@
This work highlights the need to explore alternative grouping principles and similarity measures beyond proximity and Euclidean distance to the median—particularly for more effective anomaly detection within clusters using Expectations. Broadening these foundations could enhance the robustness and adaptability of the clustering approach. As the semi-supervised algorithm relies on a small set of labelled data to initialise and guide the clustering process, future work will also investigate methods for automatically generating reliable seed points, with the goal of achieving full automation. While the algorithm’s speed and design make it well-suited for online clustering, performance profiling has identified key bottlenecks, including excessive iterations during the clustering and anomaly ejection phases, as well as areas for optimisation in the underlying anomaly detection mechanism. Future work will address the performance isses of the algorithm and implementation.